OpenStudy (lena772):

A company produces accessories for smart phones and tablets. The profit on each smart phone case is $2 and the profit on each tablet case is $3. The company made a profit of $1,200 on the cases last month. The equation 2x + 3y = 1,200 represents the company's profit from cases last month, where x is the number of smart phone cases sold and y is the number of tablet cases sold. Write the equation in function notation. Explain what the graph of the function represents. Be sure to use complete sentences.

OpenStudy (shamil98):

In function notation simply means as f(x) = so you use the slope-intercept form from the previous question.

OpenStudy (shamil98):

remember f(x) = is y =

OpenStudy (lena772):

f(x)=2?

OpenStudy (shamil98):

if y = mx + b 2x + 3y = 1200 in slope intercept form is?

OpenStudy (lena772):

y=-2/3x+400

OpenStudy (shamil98):

yes, f(x) = -2/3x + 400

OpenStudy (lena772):

Ok so how do I explain how I know that?

OpenStudy (shamil98):

hmm, it says to explain the graph.

OpenStudy (lena772):

Graph of function represents it's slope?

OpenStudy (shamil98):

The graph would represent the profit of the company. As it says so in the paragraph, if the equation represents the profit the graph would just illustrate that.

OpenStudy (lena772):

Thank you!

OpenStudy (anonymous):

so how would u do the function notation @shamil98

OpenStudy (anonymous):

@rileymathewr03 The form is f(x)=mx+b. mx is -2/3x and b is 400. so the answer is f(x)=-2/3x+400